OpenStudy (anonymous):

Evaluate the function at the given value of the independent variable and simplify. f(x) = x^2 - 1; f(x - 2)

OpenStudy (nincompoop):

okay that is easy

OpenStudy (anonymous):

I got x^2+4

OpenStudy (nincompoop):

given the function f(x) = x^2 - 1 when this function becomes f(x-2) then it would be processed as ((x-2)^2 -1)

OpenStudy (anonymous):

i dunno......i no x^2+1 isn't an option

OpenStudy (nincompoop):

x^2 -4x + 4 - 1 = x^2 -4x +3

OpenStudy (nincompoop):

I hope you understand and learned something new